OpenStudy (anonymous):

Find a quadratic equation with integral coefficients having the given roots i√ 5 , -i√ 5

OpenStudy (anonymous):

\[(x-i \sqrt{5})(x+i \sqrt{5}) = 0\] multiply it out for your quadratic equation [simply subtract the given roots]

OpenStudy (anonymous):

it would be (x+i sqrt(5))(x-i sqrt(5))=0 =x^2-(i^2)5=0 x^2+5=0

OpenStudy (anonymous):

thank you